The probability of flu symptoms for a person not receiving any treatment is

0.0440.044.

In a clinical trial of a common drug used to lower​ cholesterol,

5353

of

11201120

people treated experienced flu symptoms. Assuming the drug has no effect on the likelihood of flu​ symptoms, estimate the probability that at least

5353

people experience flu symptoms. What do these results suggest about flu symptoms as an adverse reaction to the​ drug?

​(a)

​P(Xgreater than or equals≥5353​)equals=nothing

​(Round to four decimal places as​ needed.)

​(b) What does the result from part​ (a) suggest?

A.

The drug increases the likelihood of flu symptoms because

xgreater than or equals≥5353

is highly unlikely.

B.

The drug has no effect on flu symptoms because

xgreater than or equals≥5353

is not highly unlikely.

C.

The drug increases the likelihood of flu symptoms because

xgreater than or equals≥5353

is not highly unlikely.

D.

The drug has no effect on flu symptoms because  

xgreater than or equals≥5353

is highly unlikely.
Answer #1

Ans:

a)

sample proportion=53/1120=0.047321

z=(0.047321-0.044)/SQRT(0.044*(1-0.044)/1120)

z=0.542

P(x>=53)=P(z>=0.542)=0.2939

b)correct option is B.

The drug has no effect on flu symptoms because x>=53 is not highly unlikely.